OpenStudy (anonymous):

If a,b,c are roots of the equation (x-2)(x^2 +6x-11)= 0. what is (a+b+c)

OpenStudy (sriram):

23

OpenStudy (anonymous):

i guess you have a choice. for one thing you can find the roots and just add them probably the easiest way

OpenStudy (anonymous):

it cant be 23

OpenStudy (anonymous):

it isn't it is -4

OpenStudy (sriram):

coz in ax^3+bx^2+cx+d sum of roots=-b/a

OpenStudy (anonymous):

ya it is -4

OpenStudy (anonymous):

wolframmed??

OpenStudy (sriram):

oh sorry calculation errors !!!!

OpenStudy (anonymous):

wow site just went nuts roots are 2 and \[-3\pm2\sqrt{5}\]

OpenStudy (anonymous):

but sriram is right it is \[-\frac{a_2}{a_3}=-\frac{4}{1}=-4\]

OpenStudy (anonymous):

that is why i was trying to say "you have a choice" but suddenly this site went haywire. you can multiply out as well and get \[x^3+4 x^2-23 x+22 = 0\]

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

still flashing wow yw

OpenStudy (phi):

Another way to do this is recognize one root is +2 from (x-2) the quadratic can be factored, but we know it will be 2 complex conjugate roots, so we just need the -b/2a twice, or -b/2 for the sum of the other 2 roots.